An object is shot straight up into the air from ground level with initial speed of 24 ft/sec. The height of the object (in feet) is given by the equation h=-16tsquared+24t, where t is the time in seconds after launch. Find the time(s) when the object is at ground level.
Answer by stanbon(57395)

An object is shot straight up into the air from ground level with initial speed of 24 ft/sec. The height of the object (in feet) is given by the equation h=-16tsquared+24t, where t is the time in seconds after launch. Find the time(s) when the object is at ground level.
---------
At ground level the height is zero.
To get the time let h=0 and solve for t:
-16t^2+24t=0
-8(2t^2-3t)=0
2t^2-3t=0
t(2t-3)=0
t=0 or t=3/2
This means the object was on the ground to start with (t=0)
and will be on the ground again in 1.5 seconds (t=3/2)
==========
Cheers,
Stan H.